OpenStudy (anonymous):

Determine whether each pair of equations represents parallel lines. y=2x-5,4x+2y=9 Please show work.

OpenStudy (anonymous):

no because they have different slope on is 2 and the other is negitive 2

OpenStudy (anonymous):

Writing each equation in the form \(y = mx + c\), where \(m\) is the slope 1. \(y = 2x - 5\) 2. \(y = -2x + \frac{9}{2}\), We can see the slopes are different, so the lines are not parallel.